# Calibrated Probabilistic Interpolation for GEDI Biomass

### v1

@@ 8 Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
@@ Appendix B Implementation Details
For Random Forest, we estimated epistemic uncertainty using the variance of predictions across individual trees in the ensemble. The predicted standard deviation $\sigma_{pred}$ for a given input $x$ is calculated as:
@@ Appendix E Regression Kriging
The primary limitation of regression kriging is its computational inefficiency ($O(N^{3})$). With a mean training time of over 400 seconds for a single 1x1 degree area and high variance (std 302s) depending on data density, the method is approximately 3 to 6 times slower than the ANP. The cost of variogram fitting and kriging scales poorly with the number of data points, rendering this approach intractable for the large area, high resolution mapping tasks that are the focus of this work. For this reason, we did not include it as a standard baseline in the subsequent experiments of the main body.

### v2

@@ 7 Data Availability
8 Author Contributions
RY: Conceptualization; Data curation; Formal analysis; Investigation; Methodology; Validation; Visualization; Software; Writing – original draft; Writing – review and editing.
SK: Conceptualization; Supervision; Writing – review and editing.
@@ 9 Funding
This work was supported by the Taiwan Cambridge Scholarship from the Cambridge Trust and by funding from Dr. Robert Sansom.
@@ Appendix B Implementation Details
For Random Forest, we estimated uncertainty using the variance of predictions across individual trees in the ensemble. The predicted standard deviation $\sigma_{pred}$ for a given input $x$ is calculated as:
@@ Appendix E Regression Kriging
The primary limitation of regression kriging is its computational inefficiency ($O(N^{3})$). With a mean training time of over 400 seconds for a single 1x1 degree area and high variance (std 302s) depending on data density, the method is approximately 3 to 6 times slower than the ANP. The cost of variogram fitting and kriging scales poorly with the number of data points, rendering this approach intractable for the large area, high resolution mapping tasks that are the focus of this work. Beyond just $O(N^{3})$ training complexity. Geostatistical methods require solving an optimization problem at inference time by fitting variograms, inverting covariance matrices, and computing kriging weights for each new prediction location. Even modern sparse GP approximations with inducing points only reduce it to $O(NM^{2})$ per prediction, but that is still meaningful computation that must be repeated across tens of millions of pixels for wall-to-wall mapping.
ANPs amortize this cost entirely. Training learns a function that directly maps (context set, target location) to predictive distributions. Inference is a single forward pass with fixed computational cost regardless of context size. For continental-scale mapping where predictions are needed at 10m resolution, this distinction between fast training, slow inference (sparse GPs) and moderate training, essentially free inference (ANPs) becomes operationally decisive. For this reason, we did not include it as a standard baseline in the subsequent experiments of the main body.